OpenStudy (anonymous):

Find the directional derivatives of the following functions at the given points P in the direction of the vectors v. f(x,y,z) =√xyz , P (3,2,6) , v =<-1,-2,2>

OpenStudy (anonymous):

find the derivatives of the function x, y and z seperately, by keeping the other 2 variables constant in each case. Think of a smart way to keep the other 2 variables constant!

OpenStudy (anonymous):

ok, i tried to work out. is the ans -1?

OpenStudy (turingtest):

what is you function?\[f(x,y,z)=\sqrt{xyz}\]or\[f(x,y,z)=\sqrt xyz\]?

OpenStudy (anonymous):

the first one

OpenStudy (turingtest):

\[D_{\vec v}=\nabla f(3,2,6)\cdot\frac{\vec v}{\|\vec v\|}\]I don't think the answer is one but I haven't done it yet...

OpenStudy (anonymous):

is there an advanced equation editor available to super users that has cool features like gradient?

OpenStudy (anonymous):

negative one

OpenStudy (anonymous):

@mathstina @TuringTest is right

OpenStudy (anonymous):

@mathstina first fing the derivative of the function with respect to x, by substituting the y and x values

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

oh yeah, I do get -1 :)

OpenStudy (anonymous):

ok thanks a lot!

OpenStudy (anonymous):

@TuringTest @mathstina no its not -1, atleast as per my calculations

OpenStudy (turingtest):

I will show my work and you can spot a mistake if you see one...

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

\[\nabla f=\frac12\langle\sqrt{\frac{yz}x},\sqrt{\frac{xz}y},\sqrt{\frac{xy}z}\rangle\implies\nabla f(3,2,6)=\frac12\langle2,3,1\rangle\]\[\frac{\vec v}{\|\vec v\|}=\frac13\langle-1,-2,2\rangle\]\[\nabla f(3,2,6)\cdot\frac{\vec v}{\|\vec v\|}=\frac16(-2-6+2)=-1\]

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

@TuringTest @mathstina im sorry, your right, i made a mistake by putting all -ves as +ves and vice versa. Bravo you two

OpenStudy (turingtest):

cheers!

OpenStudy (anonymous):

Its ok. u too got the same ans

OpenStudy (anonymous):

i got it as 1 previously, now i have -1 :)